### v1

@@ 3. Energy identical-relation:
$D_{\tau}E^{n+1}=-\mu\int_{\Omega}|\nabla\tilde{u}_{h}^{n+1}|^{2}\,\mathrm{d}\mathbf{x}+\int_{\Omega}f^{n+1}\tilde{u}_{h}^{n+1}\,\mathrm{d}\mathbf{x},$
where the energy $E^{n+1}$ is defined by:
$E^{n+1}=\frac{1}{2}||\sigma_{h}^{n+1}u_{h}^{n+1}||_{L^{2}}^{2}.$
@@
The positivity of $\rho_{h}^{n+1}$ can be easily derived by combining the induction method with (3.6)-(3.7).
@@ 4 Error estimate
Moreover, setting $r=r_{h}\in W_{h}$ and $(v,q)=(v_{h},q_{h})\in(V_{h},M_{h})$ in (4.2) and (4.3), subtracting (3.1) and (3.2) from (4.2) and (4.3), respectively, we have the error equations
@@
$=2\tau|(\sigma_{h}^{n+1}e_{\sigma h}^{n+1}D_{\tau}u^{n+1},\tilde{\theta}_{uh}^{n+1})|\!+\!2\tau|(\sigma_{h}^{n+1}u^{n}D_{\tau}e_{\sigma h}^{n+1},\tilde{\theta}_{uh}^{n+1})|$
$=2\tau|(\sigma_{h}^{n+1}e_{\sigma h}^{n+1}D_{\tau}u^{n+1},\tilde{\theta}_{uh}^{n+1})|\!+\!2\tau|(\sigma_{h}^{n+1}u^{n}D_{\tau}\eta_{\sigma h}^{n+1},\tilde{\theta}_{uh}^{n+1})|$
@@
Under Assumptions 2.1, there exists $\tau_{1}>0$, if $\tau<\tau_{1}$, then it is valid for the error equations (4.7), for all $10\leq n\leq N-$ 1, that
@@
$\leq C\tau^{3}||u_{h}^{n}||_{L^{\infty}}+C\tau||\sigma_{h}^{n+1}\tilde{\theta}_{uh}^{n+1}||_{L^{2}}^{2}$
@@
Finally, by utilizing (2.10), we arrive at
@@
(I-2) When $h$ and $\tau$ is sufficiently small such that $||\bar{e}_{\rho h}^{1}||_{L^{2}}^{2}\leq\epsilon_{1}<1$ with $\epsilon_{1}$ being a positive constant (see (4.46)), (4.34) in Lemma 4.5 and (4.1) imply that
@@
On the other hand, applying Theorem 2.2, we obtain
$||D_{\tau}(\Pi_{h}\sigma^{1})||_{L^{3}}$ $\leq||D_{\tau}\sigma^{1}||_{L^{3}}\!+\!||D_{\tau}\eta_{\sigma h}^{1}||_{L^{3}}\leq C\!+\!Ch||D_{\tau}\sigma^{1}||_{W^{1,3}}\!\leq C,$
@@
Taking a summation on both sides of (4.49) and applying the Gronwall inequality (2.6), we obtain
$||\sigma_{h}^{1}\tilde{\theta}_{uh}^{1}||_{L^{2}}^{2}+||\sigma_{h}^{1}\tilde{\theta}_{uh}^{1}-\sigma_{h}^{0}\tilde{\theta}_{uh}^{0}||_{L^{2}}^{2}+\mu\tau||\nabla\tilde{\theta}_{uh}^{1}||_{L^{2}}^{2}\leq C(\tau^{2}+h^{4}),$
@@
$||\theta_{uh}^{1}||_{L^{2}}^{2}$ $=|(|R_{h}u^{1}-\tilde{u}_{h}^{1})+(\tilde{u}_{h}^{1}-u_{h}^{1})||_{L^{2}}^{2}$ (4.53)
@@
Since $0\leq 1-Ch^{2}\leq\gamma_{h}^{1}\leq 1+Ch^{2}$ from (4.52), it follows that $1+\sqrt{\gamma_{h}^{1}}$ is bounded and
$|1-\sqrt{\gamma_{h}^{1}}|^{2}\leq|\frac{1-\gamma_{h}^{1}}{1+\sqrt{\gamma_{h}^{1}}}|^{2}\leq C(\tau^{2}+h^{4}).$ (4.54)
@@
(II) Assuming that (4.37) to (4.42) are valid for $k=0,1,2,\dots,m-1(1\leq m\leq N)$, following the similar process in (I), we can prove that they hold for $k=m$, too. The proof is completed.
@@ 5.2 Property-preserving test
Then, we test the property-preserving of the proposed scheme. Set $f=0,T=10$, the domain $\Omega=\left(-1,1\right)^{2}$ and the initial condition
$\rho_{0}$ $=1,$
$u_{0}$ $=(\sin(\pi x)\cos(\pi y),-\cos(\pi x)\sin(\pi y))^{\top},$
$p_{0}$ $=\frac{3}{16}\cos(2\pi x)\cos(2\pi y).$
We simulate the evolution of the density, mass, and energy for different viscosities ($\mu=0.1,0.05,0.01,0.005$). From Figure 1, we can see that the numerical density remains positive, the numerical mass is always conserved and the numerical energy is dissipative in the whole tested time interval, which is consistent with the theoretical prediction deduced above.
@@ 5.3 Back-step flow
In this section, we apply the proposed scheme to the back-step flow. With the boundary condition set in Figure 2, taking $\rho_{0}=1$, $u_{0}(\mathbf{x})=0,$ $\mu=0.01$ and $\tau=0.01$, we show the simulation results in Figures 3-5. From the results we can see that, as the time develops, the vortex appears and becomes more and more larger near the step, which is good agreement with that in the references [12].
@@ 5.4 Flow around a circular cylinder
Finally, we apply the proposed finite element scheme to the flow around a circular cylinder in this section. The domain is defined as $\Omega\in(0,6)\times(0,1)$ with no-slip boundary conditions being imposed to the top and the bottom of the channel as well as the surface of the cylinder, a circle with the radius being $0.15$ centers at $(x,y)=(1,0.5)$, and the initial velocity $u(\mathbf{x})=0$. For the simulation parameters, we set $\mu=\frac{1}{300},\tau=0.01,\rho_{0}=1,\rho|_{inflow}=1$, and the inflow boundary condition is prescribed as $u_{1}(\mathbf{x},t)=6y(1-y),u_{2}(\mathbf{x},t)=0$. While we impose the condition $-pI+\frac{\partial u}{\partial n}=0$ on the outlet. The contour plots for the velocity components $u_{1},u_{2}$ and the pressure $p$ are presented in Figures 6-8. At the beginning, both velocity and pressure are almost symmetric with respect to the line $y=0.5$ (when $t=3$). But as the time develops, the turbulence will appear (when $t=5$) and get obviously (when $t=7$) after the flow past through the circle. But their values keep symmetric with respect to the line $y=0.5$ before the circle. These are similar to that in [36]. All of these confirm the efficiency of the proposed scheme.

### v2

@@ 3. Energy identical-relation:
$D_{\tau}E_{h}^{n+1}=-\mu\int_{\Omega}|\nabla\tilde{u}_{h}^{n+1}|^{2}\,\mathrm{d}\mathbf{x}+\int_{\Omega}f^{n+1}\tilde{u}_{h}^{n+1}\,\mathrm{d}\mathbf{x},$
where the energy $E_{h}^{n+1}$ is defined by:
$E_{h}^{n+1}=\frac{1}{2}||\sigma_{h}^{n+1}u_{h}^{n+1}||_{L^{2}}^{2}.$
@@
If $\sigma_{h}^{n+1}=0$, noting that
$(\nabla\sigma_{h}^{n+1}\cdot u_{h}^{n},r_{h})+\frac{1}{2}(\sigma_{h}^{n+1}\nabla\cdot u_{h}^{n},r_{h})=$ $(\nabla\cdot(\sigma_{h}^{n+1}u_{h}^{n}),r_{h})-\frac{1}{2}(\sigma_{h}^{n+1}\nabla\cdot u_{h}^{n},r_{h})$
$=$ $-(\sigma_{h}^{n+1}u_{h}^{n},\nabla r_{h})-\frac{1}{2}(\sigma_{h}^{n+1}\nabla\cdot u_{h}^{n},r_{h}),$
substituting this equation into (3.1), we can derive that $\sigma_{h}^{n}=0$, which follows by $\sigma_{h}^{0}=\Pi_{h}\sigma^{0}=0$. It is contradictory with (2.16). Therefore, $\sigma_{h}^{n+1}\neq 0$ for all $10\leq n\leq N-$ 1, i.e., $\bar{\rho}_{h}^{n+1}>0$ by using (3.6). Then, the positivity of $\rho_{h}^{n+1}$ can be easily derived by combining the induction method with (3.6)-(3.7).
@@
Remark 3.1.
Although the energy identical-relation was considered in [44], their energy is a modified one based on the scalar auxiliary variable method, and their scheme doesn’t preserve the positivity of the density. Moreover, if the density $\rho$ is a constant, the equations (1.1)-(1.3) reduce to the classical Navier-Stokes equations, and the energy identical-relation derived in Theorem 3.1 holds in this case, too. Different from the energy dissipation law which has been widely investigated for the discrete scheme of the classical Navier-Stokes equations by assuming that the body force $f=0$ (see, i.e., [20, 23, 37]), the energy law proved here for the scheme (3.1)-(3.7) is an equality, which is a discrete analogue of the continuous property presented in Section 2.3. If $-\mu\int_{\Omega}|\nabla\tilde{u}_{h}^{n+1}|^{2}\,\mathrm{d}\mathbf{x}+\int_{\Omega}f^{n+1}\tilde{u}_{h}^{n+1}\,\mathrm{d}\mathbf{x}\leq 0$ (the body force $f=0$ can be seen as a special case under this condition), the energy of the scheme (3.1)-(3.7) will obey the dissipation law. Otherwise, the energy of the scheme (3.1)-(3.7) will increase, which means that the energy from the external body force $f$ is greater than the dissipation part of the system. This is consistent with the continuous property. The numerical example shown in Section 5 will confirm this fact.
@@ 4 Error estimate
Moreover, setting $r=r_{h}\in W_{h}\subset W$ and $(v,q)=(v_{h},q_{h})\in(V_{h},M_{h})\subset(V,M)$ in (4.2) and (4.3), subtracting (3.1) and (3.2) from (4.2) and (4.3), respectively, we have the error equations
@@
$\leq 2\tau|(\sigma_{h}^{n+1}e_{\sigma h}^{n+1}D_{\tau}u^{n+1},\tilde{\theta}_{uh}^{n+1})|\!+\!2\tau|(\sigma_{h}^{n+1}u^{n}D_{\tau}e_{\sigma h}^{n+1},\tilde{\theta}_{uh}^{n+1})|$
$\leq 2\tau|(\sigma_{h}^{n+1}e_{\sigma h}^{n+1}D_{\tau}u^{n+1},\tilde{\theta}_{uh}^{n+1})|\!+\!2\tau|(\sigma_{h}^{n+1}u^{n}D_{\tau}\eta_{\sigma h}^{n+1},\tilde{\theta}_{uh}^{n+1})|$
@@
Under Assumptions 2.1, it is valid for the error equations (4.7), for all $10\leq n\leq N-$ 1, that
@@
$\leq C\tau^{3}||u_{h}^{n}||_{L^{\infty}}+C\tau^{3}+C\tau||\tilde{\theta}_{uh}^{n+1}||_{L^{2}}^{2}+C\tau||\sigma_{h}^{n+1}\tilde{\theta}_{uh}^{n+1}||_{L^{2}}^{2}$
@@
Furthermore, by utilizing (2.10), we arrive at
@@
Finally, by employing (4.4), we obtain that
$2\tau(R_{u1}^{n+1},\tilde{\theta}_{uh}^{n+1})=C\tau||R_{u1}^{n+1}||_{L^{2}}^{2}+C\tau||\tilde{\theta}_{uh}^{n+1}||_{L^{2}}^{2}\leq C\tau^{3}+C\tau||\tilde{\theta}_{uh}^{n+1}||_{L^{2}}^{2}.$ (4.34)
@@
(I-2) There exists $\tau_{2}>0$ and $h_{0}>0$, if $\tau\leq\min\{\tau_{1},\tau_{2}\}$ and $h\leq h_{0}$, then $||\bar{e}_{\rho h}^{1}||_{L^{2}}^{2}\leq\epsilon_{1}<1$ with $\epsilon_{1}$ being a positive constant (see (4.47)), (4.35) in Lemma 4.5 and (4.1) imply that
@@
On the other hand, we can easily obtain
@@
There exists $\tau_{3}>0$, if $\tau\leq\tau*:=\min\{\tau_{1},\tau_{2},\tau_{3}\}$, then $01-C\tau>$ 0, thus we obtain
$||\sigma_{h}^{1}\tilde{\theta}_{uh}^{1}||_{L^{2}}^{2}+||\sigma_{h}^{1}\tilde{\theta}_{uh}^{1}-\sigma_{h}^{0}\tilde{\theta}_{uh}^{0}||_{L^{2}}^{2}+\mu\tau||\nabla\tilde{\theta}_{uh}^{1}||_{L^{2}}^{2}\leq C\tau^{3}+C\tau h^{4}\leq C(\tau^{2}+h^{4}),$
@@
$||\theta_{uh}^{1}||_{L^{2}}^{2}$ $=||(R_{h}u^{1}-\tilde{u}_{h}^{1})+(\tilde{u}_{h}^{1}-u_{h}^{1})||_{L^{2}}^{2}$ (4.54)
@@
Since $\tau\leq Ch^{2}$, then when h is sufficiently small, we can get $0\leq 1-Ch^{2}\leq\gamma_{h}^{1}\leq 1+Ch^{2}$ from (4.53), it follows that $1+\sqrt{\gamma_{h}^{1}}$ is bounded and
$|1-\sqrt{\gamma_{h}^{1}}|^{2}\leq\left|\frac{1-\gamma_{h}^{1}}{1+\sqrt{\gamma_{h}^{1}}}\right|^{2}\leq C(\tau^{2}+h^{4}).$ (4.55)
@@
(II) Assuming that (4.38) to (4.43) are valid for $k=0,1,2,\dots,m-1(1\leq m\leq N)$, following the similar process in (I), we can prove that they hold for $k=m$, too. The proof is completed. ∎
@@ 5.2 Property-preserving test
In this part, we test the property-preserving of the proposed scheme through two examples, which includes evolutions of the density, mass, energy and differences in the energy identical-relation with the body force $f=0$ and $f\neq 0$, respectively.
Define the differences between two sides of the energy identical-relation in Theorem 3.1 as
$D_{E}^{n}=\left|E_{h}^{n+1}-E_{h}^{n}+\mu\tau\int_{\Omega}|\nabla\tilde{u}_{h}^{n+1}|^{2}\,\mathrm{d}\mathbf{x}-\tau\int_{\Omega}f^{n+1}\tilde{u}_{h}^{n+1}\,\mathrm{d}\mathbf{x}\right|.$
Setting the time step $\tau=0.01$, the mesh size $h=0.05$, the finial time $T=50$ and the domain $\Omega=\left(0,1\right)^{2}$ with homogenous Dirichlet boundary conditions on $\partial\Omega$, we firstly test the case with the body force $f=0$ and the initial data $\rho_{0}=1,u_{0}=(10x^{2}(x-1)^{2}y(y-1)(2y-1),-10x(x-1)(2x-1)y^{2}(y-1)^{2})^{\top}$. It is easy to check that $u_{0}$ satisfies the homogenous Dirichlet boundary conditions and $\nabla\cdot u_{0}=0$. The evolutions of the density, mass, energy and $D_{E}^{n}$ for different viscosities ($\mu=0.05,0.01,0.005,0.001$) are shown in Figure 1, from which we can see that the density remains positive, the mass is always conserved and the energy is dissipative. Moreover, we can see that the differences $D_{E}^{n}$ between two sides of the energy identical-relation are close to 00. These suggests that the properties are preserved very well, which is consistent with the theoretical prediction deduced above.
Then, with the same computational environment as that in the above but replacing the body force with $f=((2+x+y)\cos(t),(2+x+y)\sin(t))^{\top}$, we investigate the evolution of density, mass, energy and $D_{E}^{n}$ for various viscosities ($\mu=0.05,0.01,0.005,0.001$) again. The simulations are presented in Figure 2. Similar results as the above are obtained for the numerical density and mass, which obey the properties derived in Theorem 3.1. But the energy is not dissipative in this case, which forms a quasi-periodic evolution due to the periodic body force $f$. Another observation is that, although the differences $D_{E}^{n}$ between two sides of the energy identical-relation are also close to 00, they almost captures the varying period of the energy, which indicates that the energy identical-relation holds, too. All of these confirms the predictions derived in Theorem 3.1.
@@ 5.3 Back-step flow
In this section, we apply the proposed scheme to the back-step flow. With the boundary condition set in Figure 3, taking $\rho_{0}=1$, $u_{0}(\mathbf{x})=0,$ $\mu=0.01$ and $\tau=0.01$, we show the simulation results in Figures 4-6. From the results we can see that, as the time develops, the vortex appears and becomes more and more larger near the step, which is good agreement with that in the references [13].
@@ 5.4 Flow around a circular cylinder
Finally, we apply the proposed finite element scheme to the flow around a circular cylinder in this section. The domain is defined as $\Omega\in(0,6)\times(0,1)$ with no-slip boundary conditions being imposed to the top and the bottom of the channel as well as the surface of the cylinder, a circle with the radius being $0.15$ centers at $(x,y)=(1,0.5)$, and the initial velocity $u(\mathbf{x})=0$. For the simulation parameters, we set $\mu=\frac{1}{300},\tau=0.01,\rho_{0}=1,\rho|_{inflow}=1$, and the inflow boundary condition is prescribed as $u_{1}(\mathbf{x},t)=6y(1-y),u_{2}(\mathbf{x},t)=0$. While we impose the condition $-pI+\frac{\partial u}{\partial n}=0$ on the outlet, where $I$ is the unit matrix of $2\times 2$. The contour plots for the velocity components $u_{1},u_{2}$ and the pressure $p$ are presented in Figures 7-9. At the beginning, both velocity and pressure are almost symmetric with respect to the line $y=0.5$ (when $t=3$). But as the time develops, the turbulence will appear (when $t=5$) and get obviously (when $t=7$) after the flow past through the circle. But their values keep symmetric with respect to the line $y=0.5$ before the circle. These are similar to that in [39]. All of these confirm the efficiency of the proposed scheme.